# DIPLODOCUS I: Framework for the evaluation of relativistic transport equations with continuous forcing and discrete particle interactions

### v2

@@ IV.2 DIP Applied to Transport
Being stationary, the functions $A,B,C,D,E,F,$ and $G$ are functions of a subset of the spatial coordinates $x^{\iota}$ only. The functions $E,F,G$ form what is conventionally known as the shift vector $\beta_{\iota}=(E,F,G)$ with the function $A$ being the lapse function. The metric given by Eq.˜42 could refer to a wide range of metrics commonly applied to astrophysical settings, an example subset of which are given in Appendix˜B.
The spacetime volume element (Eq.˜42) of the metric given by Eq.˜42 has the simple form:
@@
where $X^{\alpha}=(0,\frac{X_{1}}{B},\frac{X_{2}}{C},\frac{X_{3}}{D})$, $Y^{\alpha}=(0,\frac{Y_{1}}{B},\frac{Y_{2}}{C},\frac{Y_{3}}{D})$, and $X^{\alpha}=(0,\frac{Z_{1}}{B},\frac{Z_{2}}{C},\frac{Z_{3}}{D})$ are a set of mutually orthonormal vectors, which are also orthogonal to $\bm{n}$. It is often standard to align $\bm{X},\bm{Y},\text{ and }\bm{Z}$ to the $\{x,y,z\}$ coordinate directions, however, they have here been left general as it can be useful to align them to some other set of vectors e.g. the local magnetic and electric field directions [88, 89, 90].
With $\bm{n}$ and $e_{a}^{~\alpha}$ defined, the integral form of the transport equation Eq.˜10 can be evaluated using DIP over a phase space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the rather long expressions Eqs.˜60, 61, 62, 63, 64, 66, 67, 68, 65, 69 and 71, found in Appendix˜C, whereby terms on the left hand side of Eq.˜10 are converted into fluxes at the boundaries of the sub-domain (see Fig.˜7) and the right hand side has been converted into a series of collision arrays. These equations form a numerical evolution scheme for all particle distribution functions, including external forces and collisions between particles. Such a scheme is conservative in terms of the total number of particles in the system (in the absence of emissive processes which are free to add or subtract particles from the system), as fluxes are always balanced between sub-domains, i.e. any particles leaving one sub-domain of phase space enter another.
As an example of how these fluxes are generated, consider the the specific case of Schwarzchild geometry with coordinates $x^{\alpha}=\{t,r,\theta,\psi\}$ (see Appendix˜B). In this geometry $\bm{n}=-\sqrt{1-\frac{r_{s}}{r}}\bm{\mathrm{d}}t$, where $r_{s}$ is the Schwarzchild radius, and the coordinate transformation Eq.˜44 may be given by
@@
Using these definitions for the Eulerian observer $\bm{n}$ and coordinate transformations $e_{a}^{~\alpha}$, the fluxes through phase space sub-domain boundaries and the spacetime volume element (Eqs.˜60, 61, 62, 63, 64, 66, 67, 68, 65, 69 and 71) can be evaluated analytically. For example, the flux elements $\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}$ (Eq.˜62) through boundaries of $r=\text{constant}$ are given by:
@@
LABEL:eqn:_Bp_flux_Schwarzchild is not a particularly elegant expression but importantly it, and all other flux expressions, are independent of the distribution function and any state variables of the system. Therefore all terms relating to transport, fluxes and collisions, are pre-computable as fixed arrays.
V Conclusion
@@
In this paper, the transport of the particle distribution through phase space has been described in a general manner. Continuous forces and discrete particle interactions have been included (Sections˜II and III).
@@
A novel approach of phase space discretisation of the particle distribution function, Distribution-In-Plateaux (Section˜IV), has been described that allows for the numerical evaluation of this transport in a conservative form.
Numerical implementation and microscopic testing of this approach is detailed in Paper II, with macroscopic tests to appear in Paper III.
@@
Acknowledgements.
C. N. Everett acknowledges an Science and Technologies Facilities Council (STFC) studentship ST/X508664/1. G. Cotter acknowledges support form STFC grants ST/V006355/1, ST/V001477/1 and ST/S002952/1 and from Exeter College, Oxford. The authors would also like to thank James Matthews, Marc Klinger-Plaisier, Sera Markoff (and their research group), Pedro Ferreira, and Dion Everett for their helpful insights during various phases of this work.
@@ A.1 Binary Interactions
Consider the reversible binary (scattering) interaction $\mathfrak{a}\mathfrak{b}\rightleftharpoons\mathfrak{c}\mathfrak{d}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq.˜36:
@@
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}$ is the number of worldlines created by the forward interaction $\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}$ in this reversible process, hence giving the number of particle type-$\mathfrak{c}$ (and type-$\mathfrak{d}$) worldlines started by that interaction within a volume of phase space. Using Eq.˜35 this may be written
@@
where $p^{0}_{\mathfrak{d}}$ is now a function of the other particle momenta given by energy conservation $p^{0}_{\mathfrak{d}}=p^{0}_{\mathfrak{a}}+p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}}$ and note the shorthand for the momentum space volume element $\frac{\mathrm{d}^{3}p}{p^{0}}=\frac{(p)^{2}\mathrm{d}p\mathrm{d}u\mathrm{d}\phi}{p^{0}}=\frac{\bm{\mathrm{d}}p^{1}\wedge\bm{\mathrm{d}}p^{2}\wedge\bm{\mathrm{d}}p^{3}}{p^{0}}$.
For binary collisions, the rate is typically measured using a differential cross section. For generality of reference frame, it is best to use the Lorentz invariant differential cross section $\frac{\mathrm{d}\sigma_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}}{\mathrm{d}T}$, which is a function of the the standard Mandelstam variables $S,T\text{ and }U$. This differential cross section can be related to the scattering matrix (Eq.˜33) by
@@
The internal terms of Eq.˜49 may then be collected into the expression
@@
For the reverse process $\bm{N}_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}$, Eq.˜53 can be altered and simplified by acknowledging that now both output states (now particles of type-$\mathfrak{a}$ and type-$\mathfrak{b}$) can be integrated over, leading to
@@ A.2 Binary Interactions in DIP Form
Under the assumption that the distribution function take the form granted by DIP, Eq.˜37, the gain and loss terms Eqs.˜51 and 52 may be integrated over sub-domains of the momentum space of the particles involved to give the “gain array” components:
@@
which are pre-computable by numerical methods (see Paper II) and independent of spatial coordinates, as such are valid at any point in space.
@@ A.3 Emissive Interactions
Consider an emissive interaction where one particle “decays” into two $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq.˜36:
@@
$\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}$ is the number of worldline of particles of type-$\mathfrak{c}$ created by the emissive interaction for $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. Using Eq.˜35 this may be written:
@@ A.4 Emissive Interactions in DIP Form
Under the assumptions that the distribution of the particles takes the form granted by DIP, Eq.˜37, the gain array for an emissive interaction $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$ can be generated by integrating Eq.˜57 over discrete sub-domains of momentum space:
@@ Appendix B Metrics
A metric of the form Eq.˜42, can express most standard astrophycially relevant metrics. A subset are presented below:
[…]
@@ Appendix C DIP Transport Equations
Under the assumption that the distribution function $f(\bm{x},\bm{p})$ for all particles takes the DIP form, Eq.˜37, the integral form of the transport equation Eq.˜10, for a particle of type-$\mathfrak{c}$ can be evaluated over a phase-space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the discrete DIP transport equation:
@@
where the fluxes through hypersurfaces of constant time coordinate are:
@@
$R^{a}_{~b}=\begin{pmatrix}1&0&0&0\\ 0&\sqrt{1-u^{2}}\cos\phi&\sqrt{1-u^{2}}\sin\phi&u\\ 0&-\frac{u\sqrt{1-u^{2}}\cos\phi}{p}&-\frac{u\sqrt{1-u^{2}}\sin\phi}{p}&\frac{1-u^{2}}{p}\\ 0&-\frac{\sin\phi}{p\sqrt{1-u^{2}}}&\frac{\cos\phi}{p\sqrt{1-u^{2}}}&0\\ \end{pmatrix}$, (70)
and the spacetime volume element is
$\mathcal{V}_{\alpha\beta\gamma\delta}=\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\chi_{0123}\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z.$ (71)
The fluxes and volume elements (Eqs.˜60, 61, 62, 63, 64, 66, 67, 68, 65, 69 and 71) are all integrals over phase-space sub-domains, which in most cases can be analytically integrated as the form of all the internal functions, connection coefficients, external forces etc. are known (see, for example, LABEL:eqn:_Bp_flux_Schwarzchild).

### v4

@@ 4.2 DIP Applied to Transport
Being stationary, the functions $A,B,C,D,E,F,$ and $G$ are functions of a subset of the spatial coordinates $x^{\iota}$ only. The functions $E,F,G$ form what is conventionally known as the shift vector $\beta_{\iota}=(E,F,G)$ with the function $A$ being the lapse function. The metric given by Eq. (42) could refer to a wide range of metrics commonly applied to astrophysical settings, an example subset of which are given in Appendix B.
Figure 7: Simplified schematic of how DIP applies to phase-space transport (here phase space has been reduced to two dimensions). The action of Distribution-In-Plateaux Eq. (37), is to discretise phase space into sub-domains in which the distribution function $f$ takes a single value ($f_{ij}$ here). The left hand side of the transport equation Eq. (10) $\int_{\partial Q}f\bm{\omega}$ is then an integral over the boundaries of these sub-domains, which generates the fluxes between sub-domains ($\mathcal{A}$ and $\mathcal{B}$ here). For this two dimensional example, the transport equation for $f_{22}$ takes the form $\int_{\partial Q_{22}}f\bm{\omega}=\left(h^{-}_{32}f_{32}+h^{+}_{22}f_{22}\right)\mathcal{A}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{12}f_{12}\right)\mathcal{A}^{-}+\left(h^{-}_{23}f_{23}+h^{+}_{22}f_{22}\right)\mathcal{B}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{21}f_{21}\right)\mathcal{B}^{-}$, where the $h$ terms (Eq. (38)) dictate the value of the distribution function on sub-domain boundaries.
The spacetime volume element (Eq. (42)) of the metric given by Eq. (42) has the simple form:
@@
where $X^{\alpha}=(0,\frac{X_{1}}{B},\frac{X_{2}}{C},\frac{X_{3}}{D})$, $Y^{\alpha}=(0,\frac{Y_{1}}{B},\frac{Y_{2}}{C},\frac{Y_{3}}{D})$, and $X^{\alpha}=(0,\frac{Z_{1}}{B},\frac{Z_{2}}{C},\frac{Z_{3}}{D})$ are a set of mutually orthonormal vectors, which are also orthogonal to $\bm{n}$. It is often standard to align $\bm{X},\bm{Y},\text{ and }\bm{Z}$ to the $\{x,y,z\}$ coordinate directions, however, they have here been left general as it can be useful to align them to some other set of vectors e.g. the local magnetic and electric field directions (Chael et al., 2023; Gelles et al., 2025; Tsunetoe et al., 2025).
With $\bm{n}$ and $e_{a}^{~\alpha}$ defined, the integral form of the transport equation Eq. (10) can be evaluated using DIP over a phase space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give Eqs. (65) to (75) (found in Appendix C), whereby terms on the left hand side of Eq. (10) are converted into fluxes at the boundaries of the sub-domain (see Fig. 7) and those on the right hand side have been converted into a series of collision arrays. These equations form a numerical evolution scheme for all particle distribution functions, including external forces and collisions between particles. Such a scheme is conservative in terms of the total number of particles in the system (in the absence of emissive processes which are free to add or subtract particles from the system), as fluxes are always balanced between sub-domains, i.e. any particles leaving one sub-domain of phase space enter another.
As an example of how these fluxes are generated, consider the specific case of Schwarzchild geometry with coordinates $x^{\alpha}=\{t,r,\theta,\psi\}$ (see Appendix B, Eq. (62)). In this geometry $\bm{n}=-\sqrt{1-\frac{r_{s}}{r}}\bm{\mathrm{d}}t$, where $r_{s}$ is the Schwarzchild radius, and the coordinate transformation Eq. (44) may be given by
@@
Using these definitions for the Eulerian observer $\bm{n}$ and coordinate transformations $e_{a}^{~\alpha}$, the fluxes through phase space sub-domain boundaries and the spacetime volume element (Eqs. (65) to (75)) can be evaluated analytically. For example, the flux elements $\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}$ (Eq. (68)) through boundaries of constant $r$ are given by Eq. (LABEL:eqn:_Bp_flux_Schwarzchild).
Eq. (LABEL:eqn:_Bp_flux_Schwarzchild) is not a particularly elegant expression but importantly it, and all other flux expressions, are independent of the distribution function and any state variables of the system. Therefore all terms relating to transport, fluxes and collisions, are pre-computable as fixed arrays.
{widetext}
@@
5 Conclusion
@@
In this paper, the transport of the particle distribution through phase space has been described in a general manner. Continuous forces and discrete particle interactions have been included (Sections 2 and 3).
@@
A novel approach of phase space discretisation of the particle distribution function, Distribution-In-Plateaux (Section 4), has been described that allows for the numerical evaluation of this transport in a conservative form.
[…]
@@
Acknowledgements
C. N. Everett acknowledges a Science and Technologies Facilities Council (STFC) studentship ST/X508664/1. G. Cotter acknowledges support from STFC grants ST/V006355/1, ST/V001477/1, and ST/S002952/1, and from Exeter College, Oxford. The authors would also like to thank James Matthews, Marc Klinger-Plaisier, Sera Markoff (and their research group), Pedro Ferreira, and Dion Everett for their helpful insights during various phases of this work.
@@ A.1 Binary Interactions
Consider the reversible binary (scattering) interaction $\mathfrak{a}\mathfrak{b}\rightleftharpoons\mathfrak{c}\mathfrak{d}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq. (36):
@@
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}$ is the number of worldlines created by the forward interaction $\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}$ in this reversible process, hence giving the number of particle type-$\mathfrak{c}$ (and type-$\mathfrak{d}$) worldlines started by that interaction within a volume of phase space. Using Eq. (35) this may be written
@@
where $p^{0}_{\mathfrak{d}}$ is now a function of the other particle momenta given by energy conservation $p^{0}_{\mathfrak{d}}=p^{0}_{\mathfrak{a}}+p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}}$ and note the shorthand for the momentum space volume element $\mathrm{d}^{3}p/p^{0}=(p)^{2}\mathrm{d}p\mathrm{d}u\mathrm{d}\phi/p^{0}=\bm{\mathrm{d}}p^{1}\wedge\bm{\mathrm{d}}p^{2}\wedge\bm{\mathrm{d}}p^{3}/p^{0}$.
For binary collisions, the rate is typically measured using a differential cross section. For generality of reference frame, it is best to use the Lorentz invariant differential cross section $\frac{\mathrm{d}\sigma_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}}{\mathrm{d}T}$, which is a function of the standard Mandelstam variables $S,T\text{ and }U$. This differential cross section can be related to the scattering matrix (Eq. (33)) by
@@
The internal terms of Eq. (49) may then be collected into the expression
@@
For the reverse process $\bm{N}_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}$, Eq. (53) can be altered and simplified by acknowledging that now both output states (now particles of type-$\mathfrak{a}$ and type-$\mathfrak{b}$) can be integrated over, leading to
@@ A.2 Binary Interactions in DIP Form
Under the assumption that the distribution function takes the form granted by DIP, Eq. (37), the gain and loss terms Eqs. (51) and (52) may be integrated over sub-domains of the momentum space of the particles involved to give the “gain array” components:
@@
[…]
@@ A.3 Emissive Interactions
Consider an emissive interaction where one particle “decays” into two $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq. (36):
@@
$\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}$ is the number of worldline of particles of type-$\mathfrak{c}$ created by the emissive interaction for $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. Using Eq. (35) this may be written:
@@ A.4 Emissive Interactions in DIP Form
Under the assumptions that the distribution of the particles takes the form granted by DIP, Eq. (37), the gain array for an emissive interaction $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$ can be generated by integrating Eq. (57) over discrete sub-domains of momentum space:
@@ Appendix B Metrics
A metric of the form Eq. (42), can express most standard astrophycially relevant metrics. A subset are presented below:
Cartesian Minkowski:
$x^{\alpha}=\{t,x,y,z\},~A^{2}=1,~B^{2}=1,~C^{2}=1,~D^{2}=1,~E=F=G=0.$ (59)
Spherical Minkowski:
$x^{\alpha}=\{t,r,\theta,\psi\},~A^{2}=1,~B^{2}=1,~C^{2}=r^{2},~D^{2}=r^{2}\sin^{2}\theta,~E=F=G=0.$ (60)
Cylindrical Minkowski:
$x^{\alpha}=\{t,\rho,\vartheta,z\},~A^{2}=1,~B^{2}=1,~C^{2}=\rho^{2},~D^{2}=1,~E=F=G=0.$ (61)
Schwarzschild:
$\quad x^{\alpha}=\{t,r,\theta,\psi\},~A^{2}=\left(1-\frac{r_{s}}{r}\right),~B^{2}=\left(1-\frac{r_{s}}{r}\right)^{-1},~C^{2}=r^{2},~D^{2}=r^{2}\sin^{2}\theta,~E=F=G=0.$ (62)
Kerr (Boyer-Lindquist Coordinates):
$\begin{split}x^{\alpha}&=\{t,r,\theta,\psi\},~A^{2}=\left(1-\frac{r_{s}}{r}\right),~B^{2}=\frac{\Sigma}{\Delta},~C^{2}=\Sigma,~D^{2}=\left(r^{2}+a^{2}+\frac{r_{s}ra^{2}}{\Sigma}\sin^{2}\theta\right)\sin^{2}\theta,~E=F=0,\\ &\quad\quad G=\frac{-r_{s}ra\sin^{2}\theta}{\Sigma},~\Sigma=r^{2}+a^{2}\cos^{2}\theta,~\Delta=r^{2}-rr_{s}+a^{2}.\end{split}$ (63)
@@ Appendix C DIP Transport Equations
Under the assumption that the distribution function $f(\bm{x},\bm{p})$ for all particles takes the DIP form, Eq. (37), the integral form of the transport equation Eq. (10), for a particle of type-$\mathfrak{c}$ can be evaluated over a phase-space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the discrete DIP transport equation:
@@
where the spacetime volume element is
$\mathcal{V}_{\alpha\beta\gamma\delta}=\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\chi_{0123}\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z,$ (65)
the fluxes through hypersurfaces of constant time coordinate are:
@@
$R^{a}_{~b}=\begin{pmatrix}1&0&0&0\\ 0&\sqrt{1-u^{2}}\cos\phi&\sqrt{1-u^{2}}\sin\phi&u\\ 0&-\frac{u\sqrt{1-u^{2}}\cos\phi}{p}&-\frac{u\sqrt{1-u^{2}}\sin\phi}{p}&\frac{1-u^{2}}{p}\\ 0&-\frac{\sin\phi}{p\sqrt{1-u^{2}}}&\frac{\cos\phi}{p\sqrt{1-u^{2}}}&0\\ \end{pmatrix}$. (76)
The volume elements and fluxes (Eqs. (65) to (75)) are all integrals over phase-space sub-domains, which in most cases can be analytically integrated as the form of all the internal functions, connection coefficients, external forces etc. are known (see, for example, Eq. (LABEL:eqn:_Bp_flux_Schwarzchild)).